OpenStudy (anonymous):

what is the measure of angle LPM? 95º 85º 90º 125º

OpenStudy (anonymous):



OpenStudy (anonymous):

Angle QPL is 55?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

55+55=110 180-110=70

OpenStudy (anonymous):

Thanks!In the same garden, what is the measure of angle LPM?

OpenStudy (anonymous):

Yeah. 30+82+55+70+x Add the corner together 237+x=360 Subtract that from 360 x=123 This is angle PLM 123+30=153 Add the angles you have. 180-153=27 Subtract that from 180 to get the missing one. 27 Answer

OpenStudy (anonymous):

pellet. Wait. Messed up.

OpenStudy (anonymous):

95º 85º 90º 125º

OpenStudy (anonymous):

My bad, I know. I didn't take something into account. Hold on.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

angle PML is 30 degrees right?

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so because QN is parallel to LM, we know these two alternate interior angles are congruent